On January 1, 2016, a corporation had 10,380 shares of common stock outstanding, and on June 1, it reacquired 6,000 shares. Despite a net loss for the year of $180,000, the company declared and paid cash dividends of $24,000 and $28,000 on common and preferred stock, respectively. What was the earnings per share for 2016?


(33.72)

(18.60)

(30.23)

(22.10)
Answer #1

Earnings per share = Earnings available to common shareholders / weighted average common shares

Earnings available to common shareholders = Net income - dividend for preferred stock

= 180,000+28000 = ($ 208,000)

Weighted average common shares = (10,380 shares × 5/12) + ((10,380-6000)shares × 7/12)

= 4,325 + 2,555 = 6,880 shares

Earnings per share for 2016 = ($ 208,000) / 6,880 shares = ($ 30.23)

Hence option "c" is correct. Earnings per share for 2016 = ($ 30.23)